[figure: 9.7 cm]

MAGNETO-OPTICAL ANISOTROPY IN Co/Cu MULTILAYERS

L. Uba^{a}, S. Uba^{a}, A. Ya. Perlov^{b}, A. N. Yaresko ^{b}, V. N. Antonov^{b} and R. Gontarz^{c}

aInstitute of Physics, University at Białystok, Lipowa 41, PL-15-424 Białystok, Poland

b Institute of Metal Physics, 36 Vernadskii Street, 252142 Kiev, Ukraine cInstitute of Molecular Physics, Polish Academy of Sciences, Smoluchowskiego 17,

PL-60-179 Poznań, Poland

The dc-sputtered Co/Cu multilayers (MLS) have been investigated by spectroscopic ellipsometry and Kerr effect spectroscopy in polar and longitudinal geometry in the photon energy range 0.8-5.6 eV. The off-diagonal components of the conductivity tensor derived from the optical and magneto-optical spectra measured for different samples magnetization directions show that the system cannot be described by a single offdiagonal tensor component. The aim of the work is an investigation of the nature of the observed anisotropy between the σ_{xy} and σ_{xz} components of the effective dielectric tensor of Co/Cu MLS. To clarify the microscopic origin of the effect of magneto-optical anisotropy of Co/Cu MLS ab initio spin-polarized LMTO calculations of the model Co/Cu MLS have been performed. The relation between magneto-optical anisotropy and the structure of the MLS interfacial region has been examined and discussed.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : L. Uba

Address for correspondence : dr L. Uba Instytut Fizyki, Uniwersytet w Białymstoku ul. Lipowa 41 15-424 Białystok

Email address : ubaluba@venus.uwb.edu.pl